Martinez owns machinery that cost $87,000 with accumulated depreciation of $40,000. The company sells the machinery for cash of $42,000. The journal entry to record the sale would include:

A debit to Cash of $42,000

A credit to Gain on Sale of $2,000

A debit to Accumulated Depreciation of $47,000

A credit to Machinery of $47,000

A credit to Accumulated Depreciation of $40,000
Answer #1

Book value as on date of sale=Cost-Accumulated depreciation

=87,000-40,000=$47,000

Hence since sale proceeds is lesser than book value as on date of sale;

Hence loss =(47,000-42,000)=$5000

journal entry:

Cash a/c...Dr$42,000

accumulated depreciation a/c...Dr$40,000

Loss on sale a/c...Dr$5000

To machinery $87,000

Hence the correct option is:

A debit to Cash of $42,000